### v2

# FB-RAG: Improving RAG with Forward and Backward Lookup

Abstract
Traditional Retrieval-Augmented Generation (RAG) struggles with complex queries that lack strong signals to retrieve the most relevant context, forcing a trade-off between choosing a small context that misses key information and a large context that confuses the LLM. To address this, we propose Forward-Backward RAG (FB-RAG), a new training-free framework based on a simple yet powerful forward-looking strategy. FB-RAG employs a light-weight LLM to peek into potential future generations, using evidence from multiple sampled outputs to precisely identify the most relevant context for a final, more powerful generator. This improves performance without complex finetuning or Reinforcement Learning common in prior work. Across datasets from LongBench and Bench, FB-RAG consistently delivers strong results. Further, the performance gains can be achieved with reduced latency due to a shorter, more focused prompt for the powerful generator. On EN.QA dataset, FB-RAG matches the leading baseline with over % latency reduction or achieves an % performance improvement with a % latency reduction. Our analysis finds cases where even when the forward-looking LLM fails to generate correct answers, its attempts are sufficient to guide the final model to an accurate response, demonstrating how smaller LLMs can systematically improve the performance and efficiency of larger ones.
FB-RAG: Improving RAG with Forward and Backward Lookup
Kushal Chawla, Alfy Samuel, Anoop Kumar, Daben Liu Capital One {kushal.chawla,alfy.samuel,anoop.kumar,daben.liu}@capitalone.com
1 Introduction
Retrieval-Augmented Generation (RAG) shows immense promise in reducing hallucinations and improving generation performance fan2024survey; gao2023retrieval. RAG achieves strong results on diverse Question Answering (QA) tasks borgeaud2022improving; guu2020retrieval; asai2024self, general language tasks he2021fast; khandelwalgeneralization, and across numerous downstream applications liu2023multi; wu2024coral.
In this work, we focus on the task of answering queries based on an already-provided large context. Traditional RAG efforts for this setup involve two steps zhao2024retrieval: 1) Retrieve important chunks by computing similarities with the query (based on a sparse or dense retriever and/or a reranker), 2) Feed the retrieved chunks along with the query to an LLM, which generates the answer. We refer to these approaches as backward-looking – looking back at the input query to score the context chunks. Such methods have been widely adopted in both academia and industry. However, standard methods struggle with complex queries that lack sufficient information to retrieve relevant chunks (see example in Figure 1). This challenge is difficult to manage in RAG, where retrieving too little risks missing key information and retrieving too much risks adding irrelevant content that can confuse the LLMs yu2024defense.
To address this challenge, we design Forward-Backward RAG (FB-RAG) for studying the impact of an emerging yet underexplored idea – forward-looking or peeking into the LLM’s output generations to improve retrieval. FB-RAG generates the output in three stages: I) Recall-focused Retrieval, using an off-the-shelf retriever to extract a smaller, yet sufficiently large context, II) Precision-focused Retrieval, which either only relies on forward-looking by observing reasons and answers from a light-weight LLM to evaluate the context chunks (Ours-F) or relies on both forward and backward lookup (Ours-FB), and III) Generation, prompting a powerful LLM to get the final answer. Although prior work used related ideas to improve RAG with LLM-based feedback or confidence scores zhao2024longrag; sun2022recitation; wang2024speculative; yang2023prca; jiang-etal-2023-active, these methods typically propose complex fine-tuning or Reinforcement Learning strategies, and often assume access to external web search engines or rely on LLM’s own memory which is not suitable for many domain-specific practical settings. Instead, FB-RAG is a simple and effective training-free framework based on off-the-shelf retrievers and instruction-tuned LLMs that answers questions from an already-provided large context.
To find relevant chunks with an imperfect forward-looking LLM, FB-RAG samples multiple outputs and assigns a high score to a chunk if it was used for any of them. This turns out to be powerful for improving RAG results over recent baselines across diverse tasks. We also find that gains can be achieved while reducing latency. On EN.QA dataset from Bench zhang-etal-2024-bench, one can combine a 70B parameter model for final response generation with an 8B model for forward-lookup and match the baseline performance with over % latency reduction. Further, one can get an % performance improvement with % latency reduction. Through our qualitative analysis, we find instances where even if all the sampled outputs from the smaller LLM incorrectly answer the input query, and often fail to follow our instructions properly, this still proves sufficient for the final, more powerful LLM to generate the correct response. We now summarize our contributions:
1.
We propose FB-RAG: a novel training-free framework for performing RAG with off-the-shelf instruction-tuned LLMs. FB-RAG employs a simple and effective look-ahead strategy to evaluate context chunks before selecting them for final response generation (Section 2).
2.
We comprehensively evaluate FB-RAG against recent training-free RAG and Long Context baselines on datasets from LongBench bai-etal-2024-longbench and Bench zhang-etal-2024-bench, finding that FB-RAG delivers consistent performance improvements. We further analyze key design choices in FB-RAG, such as the number of chunks retrieved and the number of samples used for forward lookup (Sections 3 and 4).
3.
We show that FB-RAG provides the flexibility to improve performance while also reducing the latency. We additionally perform qualitative analysis discussing the strengths and limitations of our approach, and provide insights for future progress in this area (Section 5).
2 Methodology
We focus on the task of question answering based on an already-provided context. Given an input query and a context , FB-RAG relies on an off-the-shelf retriever and instruction-tuned LLMs (without finetuning) to generate the output 111This general formulation encompasses several QA, summarization, and Multiple Choice Questions (MCQ) tasks - see Section 3 for the datasets considered in this work.. We assume that context is sufficient to answer the query , differentiating from some prior formulations that assume runtime access to web search engines yan2024corrective. At its core, FB-RAG relies on a look-ahead method to retrieve the most relevant context chunks from before performing the final response generation. We start by describing this method and later connect it to the overall three-stage process of FB-RAG.
2.1 Forward-Backward Retriever
We are given a query and context , with chunks in total. We use to denote the ideal output response (ground-truth answer), and to denote the context chunk that contains the information needed to generate the ideal answer . Further, we use to represent the importance score of a context chunk given a query using an off-the-shelf retriever . We use to denote the importance score of chunk under FB-RAG given a query and the full associated context . As in a typical RAG pipeline, once the importance scores are computed, we can select the highest-scoring chunks for final output generation using an LLM. Hence, our goal in this section is simply to provide a formulation for .
Prior work has reported that LLMs often get confused by the irrelevant information present in the context xu2024recomp; asai2024self. The inverted U shape for the performance observed by yu2024defense as the context size increases demonstrates this in action. Hence, one obvious objective for the retrievers is to assign high importance scores to the most relevant chunks so that one can use a small context for generation and reduce irrelevant content. This is challenging for retrievers relying solely on the information in the input query, especially when the query is non-specific and complex li-etal-2024-retrieval. To address this gap, our key idea is to look forward at the potential answer to retrieve the relevant contexts. If we had access to the oracle generation model , we could compute in the following manner:
Unfortunately, even though we are using the oracle generator , this formulation is still not sufficient. Oftentimes in QA, the answers are concise entities or even binary (yes or no), meaning that even the ideal answer might be insufficient to identify the most relevant context chunk . Hence, we also enable the oracle to generate the ideal reasoning before generating the final answer :
For a reasonable retriever , we now hypothesize:
meaning that one can reasonably expect to reach if given access to the ideal reasoning and ideal answer . Note that our assumption that there is a single chunk which contains all the relevant information to generate is not limiting; one can trivially extend the same argument to the case where the relevant information is split across multiple chunks. In such a case, we reasonably expect the most relevant chunks to be ranked higher than irrelevant chunks based on .
We now approximate the oracle with an instruction-tuned LLM L:
where and are the reasoning and answer generated by the LLM . To capture the uncertainty of the imperfect LLM , we further propose to consider the maximum over samples generated from the model:
where and are reasoning and answer in the sample respectively. Taking the maximum ensures that even if a chunk is used only in one sample, it will still receive a high score under . This is useful to capture the relevant chunks in cases where the LLM is not confident, resulting in high variance in the samples.
Equation 5 presents the complete forward-looking component of our framework. In case of an extremely noisy model , the generated reasoning sequences and corresponding answers can be inaccurate and thus, can provide a misleading signal for our purpose of ranking the context chunks. Hence, merely relying on the outputs from such a noisy model can unfairly penalize the true relevant chunk . Motivated by this, we also incorporate a backward-looking component (as a form of a regularizer) that looks at the original input query to compute the importance scores:
where and denote the backward and forward components respectively, while and refer to their corresponding weights.
The forward component relies on (reasoning, answer) samples generated by the LLM, which can be time-consuming as is. Of course, one can generate these samples in parallel, but we propose two additional simple solutions to manage this cost. First, the LLM used for this look-ahead can be selected independently from the LLM that is used to perform the final generation. In our experiments presented in Section 4, we use a relatively light-weight LLM (B parameters) for forward-lookup and a much more powerful LLM (B parameters) for the final response generation. We also present results with other light-weight LLM choices later in Section 5. Second, one can use a fast retriever to reduce the context size before utilizing the Forward-Backward procedure laid out in this Section. These remedies motivate the three-step process of FB-RAG, which we describe below.
2.2 FB-RAG Overview
We present our approach in Figure 1. FB-RAG follows a three-stage process to compute the output response : 1) Recall-focused Retrieval, 2) Precision-Focused Retrieval, and 3) Generation.
Recall-focused Retrieval: In Stage I, we employ an off-the-shelf retriever to reduce the context size from to . This is recall-focused, meaning one can select a relatively large context while still reducing the size significantly compared to . The goal here is not to perform generation with , but rather to use it for Stage II.
Precision-Focused Retrieval: In Stage II, we follow the procedure laid out in Section 2.1 using a light-weight LLM to compute . Importantly, still comes from the full input context . We use these scores to precisely select the relevant context chunks, reducing to , which is our target context to be used for generation.
Generation: Lastly, we prompt another instruction-tuned LLM to generate the final answer.
We evaluate two variants of our framework in this paper: 1) Ours-FB: Using both and as in Equation 2.1, and 2) Ours-F: Using and (ignoring the backward-component). As presented later in Sections 4 and 5, we find that Ours-F consistently outperforms Ours-FB across the board, indicating that one needs to only rely on the forward-looking component – at least for the choices for LLM considered in this work.
We make two observations about the overall performance achievable by our framework. First, the performance is not limited by since the outputs from are only used softly to score the chunks in the full context , and the final generation is still performed by a more powerful LLM . Second, the performance is also not limited by since Stage II works to improve the position of , increasing the likelihood that is picked up in the smaller context , which can make it easier for to generate an accurate answer. We provide a deeper probabilistic interpretation of our approach in Appendix A and validate these observations empirically in Section 4.
3 Experiment Design
We address the following four research questions: RQ 1) Performance: How does FB-RAG perform compared to RAG and Long Context baselines? – We evaluate FB-RAG on datasets spanning QA, Summarization, and Multiple Choice Questions (MCQ) tasks. RQ 2) Design Considerations: What is the impact of key design choices on the performance of FB-RAG? - We study the performance by varying the number of retrieved chunks, the number of samples used in Stage II, and the LLM used for forward lookup. RQ 3) Impact on Latency: How does the three-stage process of FB-RAG impact the overall latency? - We plot the performance against latency by varying the chunks and comparing our approach to a baseline. RQ 4) Qualitative Analysis: In what specific scenarios does FB-RAG improve performance and what kind of errors does the approach make? - We perform error analysis and discuss our insights for future work.
Datasets: Following prior work li-etal-2024-retrieval, we focus on tasks that are a) in English, b) real, and c) query-based. This leads to datasets from LongBench bai-etal-2024-longbench: NarrativeQA kocisky-etal-2018-narrativeqa, Qasper dasigi-etal-2021-dataset, MultiFieldQA bai-etal-2024-longbench, HotpotQA yang-etal-2018-hotpotqa, 2WikiMultihopQA ho2020constructing, MuSiQue trivedi2022musique, and QMSum zhong-etal-2021-qmsum. We also pick two datasets from Bench zhang-etal-2024-bench, namely, En.QA and EN.MC. These datasets cover diverse domains, including Wikipedia articles, meetings, narratives, and research papers, involving single and multi-hop questions. The average context lengths range from a few thousand to k words. Refer to Appendix B for more details.
Metrics: We use F1 score for QA datasets, Rouge-L F1 for summarization, and classification accuracy for the MCQ task. Our implementation is based on the code released with LongBench222https://github.com/THUDM/LongBench/tree/main.
Methods: Long Context (LC) refers to directly feeding the full context to the LLM without explicit retrieval. Vanilla denotes the typical RAG approach, which performs retrieval based on an off-the-shelf retriever before feeding the context to the LLM. We implemented two recent approaches evaluated on the considered datasets. In Order-Preserving (OP) RAG yu2024defense, the selected chunks from the retriever are first sorted in their original ordering before feeding them to the LLM. Self-Route li-etal-2024-retrieval is a look-ahead approach that relies on LLM’s ability to understand if the question is answerable from the retrieved context. It involves steps: 1) Retrieval: Based on an off-the-shelf retriever, 2) Generation: A modified generation based on the retrieved context where the LLM can choose to output ‘unanswerable’ if it finds that the retrieved context is insufficient to answer the question, and 3) Generation: Based on the full input context if the LLM outputs ‘unanswerable’ in the previous step.
For our approach, both Ours-FB and Ours-F variants use samples in Stage II obtained by combining top-p (p=) and top-k (k=) sampling. The final response generation for all methods uses Llama3.1-70B-Instruct llama31-2024. Self-Route uses the same model for both generation steps. For our approach, we use Llama3.1-8B-Instruct llama31-2024 for generating samples in Stage II. Refer to Appendix C for the prompts used, hardware details, and token limits. We evaluated retrievers: BM25 trotman2014improvements, M3Flag chen-etal-2024-m3, BGEFlag xiao2024c, and MPNet333https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1. We chose BM25 for our experiments due to its strong relative performance, simplicity, and versatility, making it suitable for our approach, which relies on LLM-generated outputs to retrieve relevant context chunks (see Appendix D.1 for a performance comparison). For chunking, we use a chunk size of words throughout.
4 Results
Method | Avg | Narr | Qasp | Mult | Hotp | 2Wiki | Musi | QMSum |
Long Context | ||||||||
Llama3.1-70B-Instruct | 49.28 | 33.42 | 50.96 | 55.63 | 64.4 | 67.18 | 48.68 | 24.68 |
Self-Route li-etal-2024-retrieval | ||||||||
Gemini-1.5-Pro | 43.33 | 28.32 | 45.23 | 51.47 | 55.18 | 62.68 | 40.66 | 19.77 |
GPT-4O | 46.83 | 31.36 | 47.99 | 53.17 | 62.14 | 70.14 | 41.69 | 21.31 |
Llama3.1-70B-Instruct; RAG - Our Impl. (1.5k) | ||||||||
Vanilla | 44.19 | 25.01 | 49.31 | 53.41 | 60.91 | 58.84 | 37.32 | 24.51 |
OP yu2024defense | 44.34 | 23.89 | 49.31 | 54.8 | 61.11 | 59.06 | 37.94 | 24.26 |
Self-Route li-etal-2024-retrieval | 47.23 | 24.04 | 48.77 | 54.34 | 64.42 | 68.23 | 46.68 | 24.14 |
Ours-FB (6k 1.5k) | 49.36 | 30.29 | 51.38 | 56.22 | 68.76 | 63.27 | 50.92 | 24.68 |
Ours-F (6k 1.5k) | 49.36 | 28.62 | 51.29 | 55.53 | 66.99 | 65.1 | 52.93 | 25.07 |
Llama3.1-70B-Instruct; RAG - Our Impl. (3k) | ||||||||
Vanilla | 47.09 | 26.99 | 50.55 | 54.67 | 65.33 | 61.06 | 46.55 | 24.48 |
OP yu2024defense | 48.03 | 26.62 | 50.71 | 56.78 | 66.28 | 64.8 | 45.91 | 25.11 |
Self-Route li-etal-2024-retrieval | 48.29 | 27.54 | 50.09 | 56.1 | 65.64 | 66.02 | 47.75 | 24.9 |
Ours-FB (6k 3k) | 50.23 | 33.22 | 50.99 | 55.99 | 66.29 | 67.42 | 53.13 | 24.56 |
Ours-F (6k 3k) | 50.31 | 32.41 | 51.05 | 56.12 | 66.79 | 67.95 | 53.7 | 24.17 |
Llama3.1-70B-Instruct; RAG - Our Impl. (6k) | ||||||||
Vanilla | 48.59 | 31.09 | 50.12 | 55.17 | 66.39 | 65.9 | 46.72 | 24.72 |
OP li-etal-2024-retrieval | 48.75 | 29.85 | 51.35 | 55.6 | 65.53 | 65.5 | 48.85 | 24.59 |
Self-Route li-etal-2024-retrieval | 48.71 | 30.52 | 50.74 | 54.67 | 66.5 | 64.12 | 49.29 | 25.13 |
Ours-FB (6k 6k) | 50.05 | 33.24 | 50.87 | 56.57 | 65.25 | 67.76 | 51.94 | 24.75 |
Ours-F (6k 6k) | 50.51 | 34.36 | 50.84 | 57.26 | 65.36 | 67.63 | 53.4 | 24.69 |
Method | EN.QA | EN.MC |
Long Context | ||
Llama3.1-70B-Instruct | 34.26 | 71.62 |
Self-Route li-etal-2024-retrieval | ||
Gemini-1.5-Pro | 37.51 | 76.86 |
GPT-4O | 34.95 | 77.29 |
Llama3.1-70B-Instruct; OP RAG yu2024defense | ||
16k | 44.43 | 84.72 |
24k | 45.45 | 88.65 |
48k | 47.25 | 85.59 |
Llama3.1-70B-Instruct; OP RAG (Our Impl.) | ||
16k | 47.87 | 81.22 |
24k | 48.27 | 85.59 |
Llama3.1-70B-Instruct; FB-RAG (Ours) | ||
Ours-FB (24k 12k) | 49.93 | 84.28 |
Ours-FB (24k 16k) | 51.68 | 85.59 |
Ours-F (24k 12k) | 50.38 | 85.59 |
Ours-F (24k 16k) | 52.24 | 86.46 |
FB-RAG outperforms Long Context and other RAG baselines on both LongBench and Bench datasets. We present the main results on LongBench datasets in Table 1. Across diverse domains and context size settings, we find that our approach exhibits consistent performance improvements over other implemented methods. Our approach achieves the best score on out of datasets, and our method Ours-F (6k 6k), which uses a 6k context output in both Stage I and II, achieves the best average performance of . We present the results for Bench datasets in Table 2. We find performance improvements on both datasets. Our approach F (24k 16k) achieves on EN.QA outperforming both the top reported results in the OP RAG paper () and the best OP RAG result found in our own implementation (). On EN.MC, our approach achieves , which beats the best achieved in our implementation of OP-RAG () but does not beat the reported best result of , potentially due to differences in the experiment design, such as the retriever and chunking methods.
Only looking forward in Stage II of FB-RAG generally performs better than averaging out Forward and Backward components. We observe that setting in Equation 2.1 (nullifying the backward-looking component in Stage II) performs better than giving equal weight to both forward and backward looking components. This indicates that when LLM-generated reasoning and answer samples are incorporated, the input query does not seem to provide any new useful information to retrieve the most relevant context chunks, and rather hurts the ranking. This essentially points to the effectiveness of the underlying LLM used for forward lookup (Llama-3.1-8B-Instruct for these reported results). In general, the 8B model is much worse than the 70B variant used for final generation ( lower average performance in our initial experiments). Often, the former even fails to follow our formatting instructions to generate the ‘Rationale:’ and ‘Answer:’ prefixes correctly. Further, we often see the answer being absent or cut off due to the model generating a long reasoning statement, leaving no room for the answer within our hard decoding token limit. However, regardless of these issues, as long as the model outputs the appropriate language relevant to answering the input question, it helps to retrieve the most relevant chunks for the final generation step by a more powerful LLM. We also experimented with different prompts for Stage II and found that some sort of reasoning or explanation provides slight gains over only using answers (Appendix D.2).
Forward-looking improves the ranking of relevant context chunks. In Figure 2 (top), we directly compare OP-RAG with our approach on EN.QA by varying the number of chunks used for final generation444We exclude Self-Route here since it relies on LC as a fallback which already performs poorer than RAG in this case.. We find that our approach at chunks (6k context) outperforms OP RAG at chunks (24k context). On EN.MC (Appendix D.3), this happens at chunks (16k context). This goes back to the discussion in Section 2.2. With forward lookup in Stage II (albeit with a less powerful LLM), our approach essentially improves the ranking of relevant context chunks, and thus, allows one to use a smaller context for final response generation. This makes it easier for the LLM to find the correct answer, leading to improved performance.
Performance improves even with one forward sample in Stage II of FB-RAG. Finally, we analyze the impact of the number of samples used in Stage II of FB-RAG on the overall performance (Appendix D.4). We find that the performance improves greatly with only one forward sample, with maximum performance at . We also note that the trend is not strictly increasing, indicating that more samples may not always add value and this parameter must be tuned empirically.
5 Discussion
Latency Considerations: FB-RAG improves performance with lower latency. The latency of FB-RAG is governed by the two LLM calls in Stage II and III (Figure 1). We approximate the overall latency by the sum of the average time taken by Llama3.1-8B-Instruct to generate output samples in Stage II (assuming parallelization) and the average time taken by Llama3.1-70B-Instruct to generate the final answer. In Figure 2 (bottom), we plot performance against latency for EN.QA, varying the number of chunks used in Stage III and comparing to OP-RAG. This is complementary to the performance curves in Figure 2 (top). As evident, we find that FB-RAG improves performance while reducing latency. It matches the best baseline performance ( F1 at s), with over 48% reduction in latency, attaining F1 at s. Further, FB-RAG shows 8% performance improvement with a 10% reduction in latency. This can be attributed to using a lightweight 8B model for forward-lookup with a large context, and the final generation with a 70B model using a much smaller context size, and is in line with previously reported inference speedups in 8B vs. 70B variants555https://openllmbenchmarks.com/index.html.
Varying the LLM used for forward lookup: We can go even more light-weight. The latency analysis above used an 8B model for forward-lookup in Stage II of FB-RAG. Even though the 8B model fails to follow instructions properly occasionally and performs much worse compared to the 70B model, it still brings performance improvements. A natural question is – ‘Can we push this further?’ In Figure 3, we compare performance by varying the LLM used for Stage II, experimenting with Llama3.2 3B and 1B instruction-tuned variants666https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/. As evident, we find that even the 3B model shows visible improvements in performance, while the 1B performs similar to the baseline. This finding attests to the strength of FB-RAG – although the 3B variant is nearly half as accurate as the 8B model, as long as it provides the relevant language in the generated reasons and answers, it helps to retrieve the relevant context chunks for the 70B model to generate accurate answers. From these observations, we argue that FB-RAG provides the knobs to improve performance while controlling latency with reasonable design choices – this includes the number of chunks for Stage II and Stage III, and the size of the forward-lookup model.
Qualitative Analysis: Analyzing complex queries where FB-RAG decisively outperforms the baselines, we make two observations. First (which is more straightforward), there are cases where the 8B model answers the query correctly in at least one of the Stage II samples, along with giving a reasonable rationale. This directly helps to pick the relevant chunks for Stage III following Equation 5. The second situation is more interesting, where the 8B model fails to answer a multihop query in all samples. However, it answers one hop correctly in at least one of the samples, which proves to be sufficient to retrieve the correct chunks for the 70B model to handle the multiple hops correctly. Take a query from MuSiQue as an example – ‘Who is the spouse of the actor who played Hannibal Smith in the A team?’, the 8B model correctly guesses ‘George Peppard’ as the actor who played Hannibal Smith, but is unable to get to the final answer ‘Sherry Boucher‘. However, simply generating the relevant language and ‘George Peppard’ helps to retrieve the right context chunks for the 70B model to produce the correct answer – This gives insight into how even a light-weight LLM in Stage II can systematically help to improve the performance, aligned with the overall results discussed earlier.
Looking at the fewer cases where FB-RAG performs worse, we find that first, some of the errors can be traced back to the evaluation metrics. When FB-RAG predicts ‘Sebastian’ instead of ‘Sebastian Cabot’ or ‘Qatari Stars League’ instead of ‘Qatar Stars League’, it hurts the F1 score it receives. – Investing in improved metrics (potentially semantic) will be valuable in the future. Second, in some cases, the error can be attributed to the ambiguity in the input query. The answer to the question ‘The Live Life Loud album’s band signed to which label?’ is temporally dependent, and FB-RAG gets penalized since it answers correctly but from a different year than what is unfairly assumed in the ground truth answer – Incorporating the temporal dimension to curate unambiguous queries will improve the dataset quality in the future. Finally, we find cases where the 70B model fails to resolve multihop queries even with a precise input context, for instance, confusing the ‘spouse’ with the ‘mother’ of an artist – Enabling LLMs to resolve complex multihop queries is still an open, challenging problem, demanding additional dedicated efforts in this area.
6 Related Work
Long Context (LC) LLMs: Context lengths have rapidly increased, with Gemini Pro team2024gemini and Meta Llama llama4-2025 using even million tokens. However, LLMs can get confused by irrelevant parts of the context, leading to known cases where RAG significantly outperforms LC xu2023retrieval; yu2024defense. In terms of latency, LC is expensive due to the quadratically increasing compute costs with input size. We follow the RAG paradigm by first retrieving the most relevant context chunks and then feeding them to an LLM with the input query for answer generation.
Retrieval Augmented Generation (RAG): RAG has emerged as a popular paradigm competing with LC, improving performance across diverse tasks with significantly lower compute costs fan2024survey. Traditional RAG is backward-looking – the context chunks are scored based on the input query using a combination of retrievers and rerankers, which further refine the selected context gao2023retrieval. Instead, FB-RAG uses forward-looking with samples generated from an LLM to select the relevant context chunks for the final answer generation. Unlike a typical reranker, Stage II of FB-RAG selects the chunks from the full context instead of (the output of Stage 1).
Numerous efforts augment RAG with trained filters yoran2023making, trained compressors xu2024recomp, and web search engines yan2024corrective to improve retrieval quality and generation. Self-RAG asai2024self trains an LLM using special reflection tokens to retrieve on demand. li2023web and jiang-etal-2023-active perform retrieval from the web based on the LLM’s look-ahead confidence scores. Speculative RAG wang2024speculative uses a smaller trained LLM to generate answer candidates, which are then verified by another LLM. LongRAG zhao2024longrag uses plug-n-play components to extract global information and factual details from context chunks which enhances the understanding from long contexts. Our setting differs in several ways: 1) We push the performance of instruction-tuned LLMs without any further training, 2) We assume no access to external web sources, and 3) We only use forward lookup from the light-weight LLM in a soft manner for selecting relevant context chunks from the entire context, with the final generation still being performed by a more powerful LLM. Two recent papers closest to our formulation are Self-Route li-etal-2024-retrieval and Order Preserving (OP) RAG yu2024defense, which we implemented ourselves and used as baselines in this work.
7 Conclusion
We proposed and evaluated FB-RAG – a new framework for RAG with LLMs. Instead of solely relying on the input query to retrieve the relevant chunks, we employed a look-ahead mechanism tightly integrated with the task at hand. This retrieves the most relevant chunks while reducing the irrelevant information in the context, resulting in superior performance. We found that FB-RAG has the potential to improve performance while simultaneously reducing latency. We performed a qualitative analysis and discussed insights to guide future work.
Limitations
The effectiveness of FB-RAG depends on the quality of the off-the-shelf retriever that is being used. In our experiments, we found BM25 to be effective. However, if the quality of the available retriever is poor for the domain in consideration, one is forced to use a larger context size for subsequent stages, which would impact the overall latency gains from the system.
In our qualitative analysis, we observed that smaller, less-capable LLMs can be used for forward-lookup in Stage II and one can eventually get accurate responses even if this small LLM is inaccurate or fails to always follow our instructions properly. However, the minimum level of the model capability (parameters, instruction-following abilities) required for forward-looking signals to be helpful remains an open question and will be important for future investigation. Ultimately, the design choices that best manage the performance-latency tradeoff will depend on the specific application and platform constraints.
Ethical Considerations
Our work was approved by the established internal review procedure. We carefully verified the licensing information associated with all the datasets and instruction-tuned LLMs used in this work, ensuring that their use was within their intended scope. All the datasets were properly anonymized before being used. We provide dataset statistics in Appendix B and refer the readers to the original dataset papers for details regarding pre-processing steps as well as the demographics of human annotators.
All datasets considered in this work were in English. Hence, it is unclear whether our findings directly translate to other languages and cultures. However, our approach is free of any such assumptions, and we encourage future work to extend it to these other scenarios.
We further note that LLMs have been known to exhibit different kinds of gender or cultural biases that can lead to discriminatory language in the generated outputs. Hence, we call for rigorous testing before any LLM-based systems are deployed. We also recommend regular monitoring after deployment to ensure that the models’ behaviors remain within their planned scope.
Appendix A Methodology
In this Section, we provide a deeper insight into how FB-RAG works to improve the overall RAG performance. This interpretation is complementary to the discussion in Section 2.2. We lay out a probabilistic formulation of the RAG process below (extending the notation used in the main paper):
where denotes all possible contexts that can be selected in the retriever stage of RAG.
The first component, , captures the retriever’s role - a conditional probability distribution over all possible contexts that can be selected from the full context conditioned on the query . A higher probability of a specific corresponds to a higher score from the retriever and a higher likelihood of it being picked up for generation.
The second component, , captures the job of the generator - the probability of generating the answer from the query and the selected context . Note that will be high for a better quality which contains the relevant context chunks and minimizes irrelevant information, and will be low for a poor quality which misses out key relevant chunks or contains a high amount of irrelevant content.
Under this formulation, when supplied with a reasonable forward-looking LLM, the procedure laid out in Section 2.1 simply works to shift the probability mass in to better quality contexts. Combined with the better performance from the generator for these better quality contexts, this holds the potential to improve the overall probability of generating the accurate answer.
Appendix B Datasets
Our experiments are based on datasets from two popular benchmarks consisting long context lengths - LongBench bai-etal-2024-longbench and Bench zhang-etal-2024-bench. QA tasks (NarrativeQA, Qasper, MultifieldQA, HotpotQA, 2WikiMultihopQA, MuSiQue, and EN.QA) take a query and a context as input, with the goal of generating a concise answer. The summarization task (QMSum) requires generating a free-form summary based on the given query and context. For the MCQ task (EN.MC), the input additionally includes a set of choices, and the task is to choose the correct choice that answers the input query based on the provided context. We present key statistics for these datasets in Table 3.
Dataset | No. of Queries | Avg Length |
LongBench bai-etal-2024-longbench | ||
NarrativeQA | 200 | 18,395 |
Qasper | 200 | 3,599 |
MultiFieldQA | 150 | 4,539 |
HotpotQA | 200 | 9,133 |
2WikiMultihopQA | 200 | 4,873 |
MuSiQue | 200 | 11,196 |
QMSum | 200 | 10,533 |
Bench zhang-etal-2024-bench | ||
EN.QA | 351 | 150,374 |
EN.MC | 229 | 142,622 |
Appendix C Experiment Design
We provide additional experimental design details in this section to promote reproducibility. We further plan to release our code on acceptance.
C.1 Prompts
We release all the prompts used in our experiments. Tables 4 and 5 list the prompts for LongBench datasets, while Table 6 presents the prompts for the two datasets from Bench. Note that for QMSum, we use the same prompt for FB-RAG Stage II as the one used for Vanilla RAG. This is because the output summary is already descriptive, unlike other datasets where answers tend to be very concise (a few words or a phrase).
C.2 Hardware Used
All the experiments presented in this paper were performed on NVIDIA A100 GPUs. We used the default inference configuration provided by Huggingface, which uses ‘device_map=auto’. We did not use any additional optimizations.
C.3 Decoding Token Limits
We set maximum limits for the number of tokens that can be generated per LLM call. For LongBench datasets, we use the limits from the code released with the benchmark777https://github.com/THUDM/LongBench/tree/main. For EN.QA and EN.MC datasets from Bench benchmark, we set the limit to , based on the ground truth distributions. When generating both reasoning and answer in Stage II of our approach, we add to the original token limit for all datasets.
Dataset | LC, Vanilla / OP RAG | Self-Route: Stage I | FB-RAG: Stage II |
NarrativeQA | You are given a story, which can be either a novel or a movie script, and a question. Answer the question as concisely as you can, using a single phrase if possible. Do not provide any explanation. Story: {context} Now, answer the question based on the story as concisely as you can, using a single phrase if possible. Do not provide any explanation. Question: {input} Answer: | You are given a story, which can be either a novel or a movie script, and a question. Answer the question as concisely as you can, using a single phrase if possible. Do not provide any explanation. If the question cannot be answered based on the information in the article, write “unanswerable”. Story: {context} Now, answer the question based on the story as concisely as you can, using a single phrase if possible. Do not provide any explanation. If the question cannot be answered based on the information in the article, write “unanswerable”. Question: {input} Answer: | You are given a story, which can be either a novel or a movie script, and a question. Answer the question as concisely as you can, using a single phrase if possible. Story: {context} Now, first provide your reasoning briefly in 2-3 sentences starting with ’Rationale:’. Then, answer the question starting with ’Answer:’ as concisely as you can, using a single phrase if possible. Question: {input} Rationale: |
Qasper | You are given a scientific article and a question. Answer the question as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ünanswerable.̈ If the question is a yes/no question, answer ÿes,̈ n̈o,̈ or ünanswerable.̈ Do not provide any explanation. Article: {context} Answer the question based on the above article as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ünanswerable.̈ If the question is a yes/no question, answer ÿes,̈ n̈o,̈ or ünanswerable.̈ Do not provide any explanation. Question: {input} Answer: | You are given a scientific article and a question. Answer the question as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ünanswerable.̈ If the question is a yes/no question, answer ÿes,̈ n̈o,̈ or ünanswerable.̈ Do not provide any explanation. Article: {context} Answer the question based on the above article as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ünanswerable.̈ If the question is a yes/no question, answer ÿes,̈ n̈o,̈ or ünanswerable.̈ Do not provide any explanation. Question: {input} Answer: | You are given a scientific article and a question. Answer the question as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ünanswerable.̈ If the question is a yes/no question, answer ÿes,̈ n̈o,̈ or ünanswerable.̈ Article: {context} Now, first provide your reasoning briefly in 2-3 sentences starting with ’Rationale:’. Then, answer the question starting with ’Answer:’ based on the above article as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ünanswerable.̈ If the question is a yes/no question, answer ÿes,̈ n̈o,̈ or ünanswerable.̈ Question: {input} Rationale: |
MultiFieldQA | Read the following text and answer briefly. {context} Now, answer the following question based on the above text, only give me the answer and do not output any other words. Question: {input} Answer: | Read the following text and answer briefly. {context} Now, answer the following question based on the above text, only give me the answer and do not output any other words. If the question cannot be answered based on the information in the article, write “unanswerable”. Question: {input} Answer: | Read the following text and answer briefly. {context} Now, first provide your reasoning briefly in 2-3 sentences starting with ’Rationale:’. Then, answer the question starting with ’Answer:’ based on the above text. Question: {input} Rationale: |
HotpotQA | Answer the question based on the given passages. Only give me the answer and do not output any other words. The following are given passages. {context} Answer the question based on the given passages. Only give me the answer and do not output any other words. Question: {input} Answer: | Answer the question based on the given passages. Only give me the answer and do not output any other words. If the question cannot be answered based on the information in the article, write “unanswerable”. The following are given passages. {context} Answer the question based on the given passages. Only give me the answer and do not output any other words. If the question cannot be answered based on the information in the article, write “unanswerable”. Question: {input} Answer: | Answer the question based on the given passages. {context} Now, first provide your reasoning briefly in 2-3 sentences starting with ’Rationale:’. Then, answer the question starting with ’Answer:’ based on the given passages. Question: {input} Rationale: |
Dataset | LC, Vanilla / OP RAG | Self-Route: Stage I | FB-RAG: Stage II |
2WikiMultihopQA | Answer the question based on the given passages. Only give me the answer and do not output any other words. The following are given passages. {context} Answer the question based on the given passages. Only give me the answer and do not output any other words. Question: {input} Answer: | Answer the question based on the given passages. Only give me the answer and do not output any other words. If the question cannot be answered based on the information in the article, write “unanswerable”. The following are given passages. {context} Answer the question based on the given passages. Only give me the answer and do not output any other words. If the question cannot be answered based on the information in the article, write “unanswerable”. Question: {input} Answer: | Answer the question based on the given passages. The following are given passages. {context} Now, first provide your reasoning briefly in 2-3 sentences starting with ’Rationale:’. Then, answer the question starting with ’Answer:’ based on the given passages. Question: {input} Rationale: |
MuSiQue | Answer the question based on the given passages. Only give me the answer and do not output any other words. The following are given passages. {context} Answer the question based on the given passages. Only give me the answer and do not output any other words. Question: {input} Answer: | Answer the question based on the given passages. Only give me the answer and do not output any other words. If the question cannot be answered based on the information in the article, write “unanswerable”. The following are given passages. {context} Answer the question based on the given passages. Only give me the answer and do not output any other words. If the question cannot be answered based on the information in the article, write “unanswerable”. Question: {input} Answer: | Answer the question based on the given passages. The following are given passages. {context} Now, first provide your reasoning briefly in 2-3 sentences starting with ’Rationale:’. Then, answer the question starting with ’Answer:’ based on the given passages. Question: {input} Rationale: |
QMSum | You are given a meeting transcript and a query containing a question or instruction. Answer the query in one or more sentences. Transcript: {context} Now, answer the query based on the above meeting transcript in one or more sentences. Query: {input} Answer: | You are given a meeting transcript and a query containing a question or instruction. Answer the query in one or more sentences. If the question cannot be answered based on the information in the article, write “unanswerable”. Transcript: {context} Now, answer the query based on the above meeting transcript in one or more sentences. If the question cannot be answered based on the information in the article, write “unanswerable”. Query: {input} Answer: | You are given a meeting transcript and a query containing a question or instruction. Answer the query in one or more sentences. Transcript: {context} Now, answer the query based on the above meeting transcript in one or more sentences. Query: {input} Answer: |
Dataset | LC, Vanilla / OP RAG | Self-Route: Stage I | FB-RAG: Stage II |
EN.QA | Read the book and answer the question. Be very concise in your answer. Book: {context} Now, answer the question based on the book. Only give me the answer and do not output any other words. Question: {input} Answer: | Read the book and answer the question. Be very concise in your answer. If the question cannot be answered based on the information in the article, write “unanswerable”. Book: context Now, answer the question based on the book. Only give me the answer and do not output any other words. If the question cannot be answered based on the information in the article, write “unanswerable”. Question: {input} Answer: | Read the book and answer the question. Be very concise in your answer. Book: {context} Now, first provide your reasoning briefly in 2-3 sentences starting with ’Rationale:’. Then, answer the question starting with ’Answer:’ as concisely as you can. Question: {input} Rationale: |
EN.MC | Read the book and answer the question. Book: {context} Now, answer the question based on the book. Only output the answer and do not output any other words. Question: {input} {all_classes} Answer: | Read the book and answer the question. If the question cannot be answered based on the information in the article, write “unanswerable”. Book: {context} Now, answer the question based on the book. Only output the answer and do not output any other words. If the question cannot be answered based on the information in the article, write “unanswerable”. Question: {input} {all_classes} Answer: | Read the book and answer the question. Book: {context} Now, first provide your reasoning briefly in 2-3 sentences starting with ’Rationale:’. Then, answer the question starting with ’Answer:’ as concisely as you can. Question: {input} {all_classes} Rationale: |
Appendix D Results
D.1 Retriever comparisons
We compared the performance of several off-the-shelf retrievers in our initial experiments, as presented in Table 7. All methods use OP RAG at 3k context size. We find that BM25 performs reasonably well on average in comparison to numerous top-performing semantic retrievers. In addition, BM25 is a versatile approach without any underlying assumptions about the query, making it well-suited for our forward-looking approach in this paper. Hence, we fixed BM25 as the retriever for the rest of our experiments discussed in Section 4 in the main paper.
D.2 FB-RAG Stage II Prompt comparisons
We experimented with a few prompt variations for Stage II of FB-RAG. Table 8 presents these comparisons on LongBench datasets. We observe that only using the answers sampled from the LLM shows improvements over other RAG baselines presented in the main paper, although this can be further improved slightly by using some form of reasoning along with the answer. This helps to handle scenarios where the answers are entity names or binary that contain little information for retrieving the most relevant context chunks.
D.3 Varying the number of chunks used for final generation
In Figure 4, we compare the performance of our approach with OP-RAG on EN.MC dataset by varying the number of chunks used for final generation. We find that FB-RAG at chunks (16k context) beats the best performance of the baseline at chunks (24k context).
D.4 Varying the number of samples used in Stage II of FB-RAG
We present the plot for analysis in Figure 5. The X-axis denotes the number of samples used. The Y-axis denotes the average performance on LongBench datasets. The results are shown for the Ours-FB (6k 3k) configuration. As evident from the figure, we find that the performance improves visibly with just one forward sample, while attaining the maximum at samples.
Method | Avg | Narr | Qasp | Mult | Hotp | 2Wiki | Musi | QMSum |
BM25 | 48.03 | 26.62 | 50.71 | 56.78 | 66.28 | 64.8 | 45.91 | 25.11 |
M3Flag (1, 0, 0) | 48.3 | 29.4 | 50.36 | 55.99 | 63.76 | 66.47 | 47.87 | 24.23 |
M3Flag (1, 0.3, 0) | 48.58 | 29.79 | 50.14 | 55.86 | 64.83 | 66.78 | 48.33 | 24.36 |
BGEFlag | 48.05 | 27.79 | 51.24 | 53.99 | 66.64 | 66.46 | 45.74 | 24.49 |
MPNet | 46.92 | 25.97 | 50.72 | 54.33 | 62.95 | 65.55 | 44.7 | 24.25 |
Method | Avg | Narr | Qasp | Mult | Hotp | 2Wiki | Musi | QMSum |
Only answers | 50.09 | 30.63 | 52.11 | 56.17 | 66.16 | 68.97 | 51.49 | 25.07 |
Thought process | 50.09 | 32.33 | 51.6 | 55.63 | 65.42 | 68.09 | 52.8 | 24.76 |
Explanation | 50.33 | 30.83 | 51.84 | 55.88 | 66.92 | 68.62 | 53.67 | 24.54 |
Reasoning | 50.23 | 33.22 | 50.99 | 55.99 | 66.29 | 67.42 | 53.13 | 24.56 |